OpenStudy (anonymous):

Explain the difference in solving for x in the following situations: x squared = 121 and x = the square root of 121.

OpenStudy (anonymous):

\[x^2=(11)^2 \therefore x=11\]

OpenStudy (anonymous):

The former has two solutions and the later one.

OpenStudy (anonymous):

\[\sqrt{x} =\sqrt{121) \therefore }\]